An experiment shows that a 114 mL gas sample has a mass of 0.173 g at a pressure of 718 Torr and a temperature of 33 ∘C.

Whats the molar mass of gas
Answer #1

Using ideal gas law:

PM = dRT

Where ,

P = pressure in atm = ( 718/760 ) atm = 0.945 atm

M = molar mass in g/mol

d = Density = (0.173 g )/(114 × 10-3 L) = 6.25 g/L

R = Gas constant = 0.0821 L•atm/mol•K

T = Temperature = 33℃ = 33+273 = 306 K

Substitute the values,

(0.945 atm)×(M) = (6.25 g/L) × (0.0821 L•atm/mol•K)×(306 K)

M = 166.154 g/mol